OpenStudy (anonymous):

Rich asked his friends how many times they went to the movies each year. Movies per Year Friend Times Tim 10 Sonya 24 Cathy 6 Martina 5 Edmundo 5 Which measure would Rich most likely use to show his parents that his friends go to the movies more than he does? Answers bubbles coming up

OpenStudy (anonymous):

A. Median. The reason: It disregards the highest and lowest amounts and would be representative of the number of times his friends go to the movies. B. Mode. The reason: It gives the number of times that two of his friends get to go to the movies. C. Mean. The reason: The mean of 24 will raise the average so it looks like everyone gets to go to more movies.

OpenStudy (anonymous):

Please help! I have few more minutes

OpenStudy (anonymous):

A. When finding the median first of all you have to list the numbers is ascending order. then find the middle value of the number list. easy. ascending order - 5 , 5 , 6 , 10 , 24 and the median is 6. B. mode is the highest frequency number that occurs. and there are 2 two fives. so the mode is 5. C. mean is the average. mean = (5 + 5 + 6 + 10 + 24)

OpenStudy (anonymous):

C. mean is the average mean = (5 + 5 + 6 + 10 + 24) / 5 = 50 / 5 = 10 earlier C answer is unfinished.

OpenStudy (anonymous):

Thanks